OpenStudy (anonymous):

if x/y=3/5 then x/3=?

OpenStudy (anonymous):

5x=3y 5x/3 = y x/3 = y/5

OpenStudy (anonymous):

:) that's right

OpenStudy (anonymous):

thank you